OpenStudy (anonymous):

what line having a slope of 3/4 passes through the point {-8, 4}

OpenStudy (anonymous):

use the equation of a line: y=mx+b

OpenStudy (anonymous):

plug in 3/4 for the slope, plug in -8 for x and 4 for y, solve for b

OpenStudy (anonymous):

y=3/4 x + 10

OpenStudy (anonymous):

(4) = (3/4)(-8) + b

OpenStudy (anonymous):

4 = -24/4 + b

OpenStudy (anonymous):

4 = -6 + b

OpenStudy (anonymous):

b = 10

OpenStudy (anonymous):

y=mx+b y = (3/4) x + b

OpenStudy (anonymous):

y = (3/4)x + 10

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

np :)